# Vertex Block Descent

### v2

@@ Abstract.
We present and evaluate our method in the context of elastic body dynamics, providing details of all essential components. Then, we discuss how it can be used for other simulation problems, including particle-based simulations and rigid bodies.
@@
ccs: Computing methodologies Physical simulationccs: Computing methodologies Collision detection
@@ 1. Introduction
In this paper, we introduce vertex block descent (VBD), a physics solver that offers unconditional stability, superior computational performance than prior methods, and the ability to achieve numerical convergence. Though our method is a general solution that can be used for a variety of simulation problems, we present and evaluate it in the context of elastic body dynamics. Then, we briefly discuss how our method can be applied to some other simulation systems, including particle-based simulations and rigid bodies.
@@ 3.2. Local System Solver
Another challenge with the analytical solver arises when encountering a (nearly) low-rank Hessian. To address this, we propose a simple solution: if $|\det(\vec{H}_{i})|\leq\epsilon$ for some small threshold $\epsilon$, we opt to bypass adjusting this particular vertex for that iteration. Given that its neighboring vertices are likely to undergo adjustments before the next iteration, it is improbable that its Hessian will remain low-rank in subsequent iterations. With this simple solution, in the extreme scenario where all vertices possess a low-rank Hessian, the system could potentially become frozen. Yet, it is crucial to note that such a case is highly improbable, since the Hessian of the inertia potential is always full-rank. One potential remedy for this would be switching to the modified Conjugate Gradient solver (Lan et al., 2023) when such a case happens. However, doing this will add additional branching to the code and can slow down the solver. Thus, we have not included it in our implementation, but, depending on specific use cases, there is always the flexibility to opt for the Conjugate Gradient solver as a backup solution.
@@ 3.4. Constraints
Since our method directly manipulates the position of each vertex, managing a constraint on a vertex becomes straightforward. Constraints generally fall into two categories: unilateral ($C(\mathbf{x})\leq 0$) or bilateral ($C(\mathbf{x})=0$). With biliteral constraints, if a vertex position is directly set to a specific value, we simply skip updating its position. Otherwise, it is constrained to a (usually linear) subspace. Our approach involves representing the constrained vertex position using the subspace basis. This transforms both the vertex position and gradient into an $L$-dimensional vector, where $L$ is the subspace dimension. Consequently, handling local steps for constrained vertices involves solving an $L\times L$ system. Regarding unilateral constraints, we allow compromises and define potential energy to be solved alongside other potentials. This method handles world box constraints in our simulations.
@@ 3.7. Initialization
Since our method is an iterative solver, we begin with an initial guess for $\vec{x}$. The closer this initial guess is to the resulting $\vec{x}^{t+1}$, the fewer number of iterations we would need to converge. Regardless of the initialization strategy for determining this initial guess, because we are solving an optimization problem, the system is expected to converge to the same result, though it may require a different number of iterations.
@@
Option (a) struggles with substantially stiff materials. As we adjust each vertex separately, assuming the others remain fixed, our method must lean on the inertia potential’s gradient (i.e. ${m_{i}(\mathbf{x}_{i}-\mathbf{y}_{i})/2h^{2}}$) to march toward the local minimum. With stiff materials, this method results in notably slower convergence rates due to the inertia potential being considerably less stiff. Consequently, it encounters challenges in simulating scenarios that resemble free fall, like a swinging elastic pendulum at its maximum height, as shown in 4(a).
Option (b) allows the system to start with the inertia of the previous step, which helps, but terminating the iterations prior to convergence can again result in local material stiffness overpowering external acceleration, as shown in 4(b).
Option (c) is similar to the initialization of position-based dynamics, and performs notably better as it effectively preserves inertia and properly reacts to external acceleration. However, with a limited number of iterations, materials behave softer than they should, often resulting in overstretching or collapsing under gravity. An example of this can be seen in 4(c), where the pendulum extends more than it should. Most notably, it struggles with steady objects at rest in contact, consistently initializing them into a penetrating state, as if they are in free fall. In such cases, the contact forces must entirely undo the position change of initialization during the iterations. This not only creates unnecessary computation, but also places considerable strain on the accuracy of the collision detection and handling methods (including friction). Therefore, properly simulating objects that are stacked on top of each other becomes a major challenge with this initialization option.
@@
This adaptive approach includes external acceleration in the initialization when the motion of a vertex resembles free fall. When an object is stationary, however, as in rest-in-contact, it maintains the previous position in the initialization, preventing undesired penetrations before the first iteration. It also successfully avoids excessive stretching, as can be seen in 4(d). Visualizations of different $\tilde{a}$ values in this simulation are shown in Figure 6. In short, our adaptive initialization is a simple but effective strategy and it is possible because VBD does not dictate a particular initialization (unlike XPBD, for example).
@@ 3.9. Parallelization
When all force elements are known ahead of the simulation, graph coloring can be performed as a preprocess. However, dynamically generated constraints/force elements, such ones due to collisions, cannot be known ahead of time, requiring dynamic recoloring.
@@
We resolve this by having an auxiliary vertex position buffer that stores the updated vertex position. When executing each local VBD position update, we write the updated vertex positions to the auxiliary buffer instead of directly overwriting the original vertex position buffer. After the position updates for each color, we copy the updated positions from the auxiliary buffer to the original vertex position buffer. This method prevents the race condition that arises from simultaneous read and write operations on a vertex position.
Under this mechanism, dynamically generated force elements using multiple vertices of the same color result in (partially) Jacobi-style iterations for those vertices. This is because the process relies on the positions from the previous iteration of vertices sharing the same color. For different-colored vertices, it is equivalent to a GS approach, considering the updated positions of other colors. Note that this differentiation is implicitly handled through our racing-free scheme, without explicitly separating the approaches in our code.
@@
3.10. GPU Implementation
In this section, we describe a GPU implementation specifically designed to leverage the inherent parallelism of modern GPUs, which consists of two hierarchical levels: block-level and thread-level parallelism. Block-level parallelism facilitates large-scale parallel operations, with the stipulation that each block operates independently. On the other hand, thread-level parallelism provides finer, single-instruction-multiple-thread (SIMT) style parallelism, allowing for inter-thread communication and synchronization within the same block.
Reflecting on Vertex Buffer Data (VBD), we observed that it naturally aligns with this hierarchical architecture. We have thousands of vertices within a single color category that operate independently, and each vertex is associated with multiple force elements, which can be processed in parallel. Hence we provide the algorithm described by Algorithm 1. The cornerstone of this strategy involves assigning a thread block to each vertex (line 7), as opposed to using a single thread. Within these blocks, each thread concurrently computes the force and Hessian for an adjacent force element, storing the outcomes in the shared memory (lines 8$\sim$11). We allocate a fixed number of threads for each thread block, denoted by $n_{t}$. When the total number of adjacent force elements surpasses $n_{t}$, individual threads are tasked with looping over their assigned elements. During this process, they calculate the forces and Hessians for these elements and then collectively sum these values in their assigned shared memory. At last, the forces and Hessians of all the threads are combined using a local reduction sum method (lines 12,13).
This parallelization mechanism turns out to be remarkably efficient. In our tests, it runs nearly an order of magnitude faster than the trivial parallelization, where a thread is assigned to each vertex to sequentially process its adjacent force elements. The primary advantage lies in the optimization of the memory access pattern, a common bottleneck in GPU programs. This method notably minimizes memory divergence within warps. Given that elements adjacent to a vertex create an n-dimensional disc (2 for triangular mesh and 3 for tetrahedral mesh), these neighboring elements often share multiple vertices. For instance, in a tetrahedral mesh, a vertex $\vec{x}_{a}$ adjacent to another vertex $\vec{x}_{b}$ is shared by at least three, but more commonly over five, neighboring tetrahedra of $\vec{x}_{b}$. This shared access can propagate to all threads within the same warp, making a single global memory access far more efficient. Approximately, in our parallelism scheme, one global memory access can be shared by about 5 threads.
Furthermore, this strategy facilitates parallel processing in evaluating the force and Hessian of adjacent elements, which are then written to the significantly faster shared memory. This bypasses the slower global memory and enables the parallel aggregation of force and Hessian values, enhancing the overall efficiency of the process.
@@ 3.10. GPU Implementation
1 $\vec{y}\leftarrow\vec{x}^{t}+h\vec{v}^{t}+h^{2}\vec{a}_{\text{ext}}$ Initial DCD using $\vec{x}^{t}$ $\vec{x}\leftarrow$ initial guess with adaptive initialization for each iteration $n\leq n_{\max}$ do
2      if $n\!\!\mod{n_{\text{col}}}=1$ then CCD using $\vec{x}$ for each color $c$ do
           // block level parallelization
@@
                // thread level parallelization
@@
5                     $\mathbf{f}_{i,j}=-\frac{\partial E_{j}}{\partial\mathbf{x}_{i}}$ $\mathbf{H}_{i,j}=\frac{\partial^{2}E_{j}}{\partial\mathbf{x}_{i}\partial\mathbf{x}_{i}}$
@@
7                     $\mathbf{f}_{i}=\sum_{j\in\mathcal{F}_{i}}\mathbf{f}_{i,j}$ $\mathbf{H}_{i}=\sum_{j\in\mathcal{F}_{i}}\mathbf{H}_{i,j}$ $\Delta\vec{x}_{i}\leftarrow\mathbf{H}_{i}^{-1}\mathbf{f}_{i}$ $\Delta\vec{x}_{i}\leftarrow$ optional line search from $\vec{x}_{i}+\Delta\vec{x}_{i}$ to $\mathbf{x}_{i}$ $\vec{x}_{i}^{\text{new}}\leftarrow\vec{x}_{i}+\Delta\vec{x}_{i}$
@@
                    // Copy updated positions back to the vertex buffer to achieve GS iteration
@@ 10                          $\mathbf{x}_{i}=\vec{x}_{i}^{\text{new}}$
$\vec{v}=(\vec{x}-\vec{x}^{t})/h$ return $\vec{x}$, $\vec{v}$
Algorithm 1 VBD simulation for one time step on GPU.
@@
                                                                   $\uparrow$ This platform will be removed.
@@ 4. Results
We use a fixed frame time of $1/60$ seconds and a fixed iteration count $n_{\max}$, instead of estimating convergence after every iteration. Each frame is computed using $S$ substeps, such that ${h=1/(60S)}$ seconds. Using smaller time steps increases accuracy and reduces numerical damping with any implicit Euler method. With VBD, a smaller time step only requires fewer iterations per step for similar visual quality, but it can also achieve a smaller residual with the same number of total iterations per frame (i.e., $Sn_{\max}$).
@@ 4.1. Large-Scale Tests
As can be seen in our supplemental video, both of these simulations exhibit stable motion, quickly forming static piles, while maintaining rest-in-contact behavior with over 1 million active collisions. VBD’s parallelism and fast convergence resulted in an average computation time of 16.4 and 7.2 seconds per frame in these simulations, respectively.
Figure 13. Visual convergence with different numbers of iterations per frame for different material stiffness (with accelerated iterations using ${\rho=0.93,0.86,0.75}$ top to bottom), simulating a beam with 463 vertices and 1.5 thousand tetrahedra.
@@ 4.4. Convergence Rate
The first alternative is preconditioned gradient descent (GD) (Wang and Yang, 2016) implemented on GPU within the same framework as ours. GD requires a form of line search for stability, which is implemented as testing the variational energy after every 8 iterations and, when needed, reducing the optimization step size and goes back to the state 8 iterations ago (following the implementation of Wang and Yang (2016)). We also include a version of GD that is accelerated using the Chebyshev semi-iterative approach (Wang, 2015), as our method. Although GD iterations are approximately 30% faster than our method, it necessitates a line search to prevent explosion at the given state. With line search, it is 10% slower than our method. Its convergence rate per iteration is considerably slower, as it corresponds to Jacobi iterations. In this example, when combined with acceleration, GD performs similarly to our method without acceleration but lags significantly behind our method with acceleration.
@@
We also compare to a version of our method that uses Jacobi iterations, called Block Jacobi, implemented by computing the position change for all vertices in parallel and then applying the position update simultaneously to all vertices at the end of each iteration. We incorporate the same line search scheme as GD for Block Jacobi, as it is necessary to prevent explosion. Block Jacobi outperforms GD, as it uses vertex blocks for computation, which corresponds to using a diagonal Hessian block as a precondition, as opposed to just a diagonal line that GD uses. Without the line search, it runs 20% faster than the GS version, due to its perfect vertex-level parallelization (without any coloring). Combined with line search, it is 17% slower than the GS version. Because it uses Jacobi iterations, its convergence is hindered, as compared to our Gauss-Seidel iterations.
@@
Furthermore, we compare our method’s convergence to Newton’s method using a direct Cholesky (LDLT) factorization solver provided by Intel’s MKL library (Wang et al., 2014). To ensure convergence, we do a PSD projection for each tet’s Hessian of elasticity. Newton’s method uses a line search for each iteration. Since its iterations are slow, the computational overhead of line search is negligible. Though the convergence of Newton’s method per iteration is far superior to all others, because of its expensive computation time per iteration, in these examples it lags far behind. Nonetheless, for relatively tame experiments with less stretching and motion, and especially for highly stiff and high-resolution simulations that are much more expensive to simulate, we would expect Newton’s method to eventually overtake all alternatives beyond a certain level of convergence.
@@
Experiment Name Number of Mateiral Contact &Fiction Simulation Parameters Time per step
@@
Squishy Ball Drops (Figure.7,9,18) 230K 700K 8 NeoHookean $\mu=2e6,\lambda=2e7$ 1e-7 1e7 0.1, 1e-2 1/120 120 15/17ms
@@
2 Cube colliding (Figure 19) 800 2.9K 6 NeoHookean $\mu=1e6,\lambda=1e7$ 1e-6 1e7 0.3, 1e-2 1/300 10 0.21/0.27ms

### v4

@@ Abstract.
We present and evaluate our method in the context of elastic body dynamics, providing details of all essential components and showing that it outperforms alternative techniques. In addition, we discuss and show examples of how our method can be used for other simulation systems, including particle-based simulations and rigid bodies.
@@
article: 116copyright: rightsretainedjournal: TOGjournalyear: 2024journalvolume: 43journalnumber: 4publicationmonth: 7doi: 10.1145/3658179ccs: Computing methodologies Physical simulationccs: Computing methodologies Collision detection
@@ 1. Introduction
In this paper, we introduce vertex block descent (VBD), a physics solver that offers unconditional stability, superior computational performance than prior methods, and the ability to achieve numerical convergence to an implicit Euler integration. Though our method is a general solution that can be used for a variety of simulation problems, we present and evaluate it in the context of elastic body dynamics. Then, we briefly discuss how our method can be applied to some other simulation systems, including particle-based simulations and rigid bodies.
@@ 3.2. Local System Solver
Another challenge with the analytical solver arises when encountering a (nearly) rank-deficient Hessian. To address this, we propose a simple solution: if $|\det(\vec{H}_{i})|\leq\epsilon$ for some small threshold $\epsilon$, we opt to bypass adjusting this particular vertex for that iteration. Given that its neighboring vertices are likely to undergo adjustments before the next iteration, it is improbable that its Hessian will remain rank-deficient in subsequent iterations. With this simple solution, in the extreme scenario where all vertices possess a rank-deficient Hessian, the system could potentially become frozen. Yet, it is crucial to note that such a case is highly improbable, since the Hessian of the inertia potential is always full-rank. One potential remedy for this would be switching to the modified Conjugate Gradient solver (Lan et al., 2023) when such a case happens. However, doing this will add additional branching to the code and can slow down the solver. Thus, we have not included it in our implementation, but, depending on specific use cases, there is always the flexibility to opt for the Conjugate Gradient solver as a backup solution.
@@ 3.4. Constraints
Since our method directly manipulates the position of each vertex, managing a constraint on a vertex becomes straightforward. Constraints generally fall into two categories: unilateral ($C(\mathbf{x})\leq 0$) or bilateral ($C(\mathbf{x})=0$). With bilateral constraints, if a vertex position is directly set to a specific value, we simply skip updating its position. Otherwise, it is constrained to a (usually linear) subspace. Our approach involves representing the constrained vertex position using the subspace basis. This transforms both the vertex position and gradient into an $L$-dimensional vector, where $L$ is the subspace dimension. Consequently, handling local steps for constrained vertices involves solving an $L\times L$ system. Regarding unilateral constraints, we allow compromises and define potential energy to be solved alongside other potentials. This method handles world box constraints in our simulations.
@@ 3.7. Initialization
Since our method is an iterative solver, we begin with an initial guess for $\vec{x}$. The closer this initial guess is to the resulting $\vec{x}^{t+1}$, the fewer number of iterations we would need to converge. Typically, if the simulation converges, different initializations should not significantly affect the final results, although they may influence the number of iterations required to achieve convergence.
@@
Option (a) struggles with substantially stiff materials. As we adjust each vertex separately, assuming the others remain fixed, our method must lean on the inertia potential’s gradient (i.e. ${m_{i}(\mathbf{x}_{i}-\mathbf{y}_{i})/2h^{2}}$) to march toward the local minimum. With stiff materials, this method results in notably slower convergence rates due to the inertia potential being considerably less stiff. Consequently, it encounters challenges in simulating scenarios that resemble free fall, like a swinging elastic pendulum at its maximum height, as shown in 5(a).
Option (b) allows the system to start with the inertia of the previous step, which helps, but terminating the iterations prior to convergence can again result in local material stiffness overpowering external acceleration, as shown in 5(b).
Option (c) is similar to the initialization of position-based dynamics, and performs notably better as it effectively preserves inertia and properly reacts to external acceleration. However, with a limited number of iterations, materials behave softer than they should, often resulting in overstretching or collapsing under gravity. An example of this can be seen in 5(c), where the pendulum extends more than it should. Most notably, it struggles with steady objects at rest in contact, consistently initializing them into a penetrating state, as if they are in free fall. In such cases, the contact forces must entirely undo the position change of initialization during the iterations. This not only creates unnecessary computation, but also places considerable strain on the accuracy of the collision detection and handling methods (including friction). Therefore, properly simulating objects that are stacked on top of each other becomes a major challenge with this initialization option.
@@
This adaptive approach includes external acceleration in the initialization when the motion of a vertex resembles free fall. When an object is stationary, however, as in rest-in-contact, it maintains the previous position in the initialization, preventing undesired penetrations before the first iteration. It also successfully avoids excessive stretching, as can be seen in 5(d). Visualizations of different $\tilde{a}$ values in this simulation are shown in Figure 6. In short, our adaptive initialization is a simple but effective strategy and it is possible because VBD does not dictate a particular initialization (unlike XPBD, for example).
@@ 3.9. Parallelization
When all force elements are known ahead of the simulation, graph coloring can be performed as a preprocess. However, dynamically generated constraints/force elements, such as ones due to collisions, cannot be known ahead of time, requiring dynamic recoloring.
@@
We resolve this by having an auxiliary vertex position buffer ($\mathbf{x}^{\text{new}}$) that stores the updated vertex position. When executing each local VBD position update, we write the updated vertex positions to the auxiliary buffer, instead of directly overwriting the original vertex position buffer. Then, we copy the updated positions from the auxiliary buffer to the original vertex position buffer after each color pass. This prevents the race conditions that arise from simultaneous read and write operations on vertex positions.
With this process, dynamically generated force elements using multiple vertices of the same color result in (partially) Jacobi-style iterations for those vertices, because they rely on the positions from the previous iteration of those vertices. For vertices with different colors, it is equivalent to Gauss-Seidel iterations, considering the updated positions of the vertices with different colors. Note that our algorithm does not explicitly switch between Jacobi and Gauss-Seidel iterations, but the resulting iteration we describe above corresponds to either (partially) Jacobi or Gauss-seidel, depending on the colors of the colliding vertices.
@@
4. GPU Implementation
In this section, we describe a GPU implementation specifically designed to leverage the inherent parallelization mechanism of modern GPUs, which consists of two hierarchical levels: block-level and thread-level parallelism. Block-level parallelism facilitates large-scale parallel operations, assuming that each block operates independently. On the other hand, thread-level parallelism provides finer, single-instruction-multiple-thread (SIMT) style parallelism, allowing for inter-thread communication and synchronization within the same block.
Reflecting on VBD, we observed that it naturally aligns with this hierarchical architecture. We have thousands of vertices within a single color category that operate independently, and each vertex is associated with multiple force elements, which can be processed concurrently. Algorithm 1 shows the pseudocode of our implementation. It uses block-level parallelism for processing each vertex. The threads within each block are used for computing the forces and Hessians, storing them in local shared memory, and computing the sums via reduction. We use a fixed number of threads for each block. When the total number of adjacent force elements exceeds the number of threads for each block, individual threads will loop over their assigned elements. During this process, they calculate the forces and Hessians for these force elements and then sum them to their assigned shared memory.
In our experiments, we observed nearly an order of magnitude performance improvement, as compared to processing each vertex with a single thread. The primary advantage lies in the optimization of the memory access pattern, a common bottleneck in GPU programs. This implementation reduces memory divergence within blocks. Because the neighboring force elements of a vertex often share multiple vertices, the threads within the same block can share a global memory access to those shared vertices. Furthermore, this strategy improved the parallelism of the algorithm by allowing parallel evaluation of the force and Hessian of adjacent force elements, which are then written to the significantly faster shared memory. This bypasses the slower global memory and enables the parallel aggregation of force and Hessian values, enhancing the overall efficiency of the process.
@@ 4. GPU Implementation
1 $\vec{y}\leftarrow\vec{x}^{t}+h\vec{v}^{t}+h^{2}\vec{a}_{\text{ext}}$
2 Initial DCD using $\vec{x}^{t}$
3 $\vec{x}\leftarrow$ initial guess with adaptive initialization
4 for each iteration $n\leq n_{\max}$ do
5      if $n\!\!\mod{n_{\text{col}}}=1$ then CCD using $\vec{x}$
6
7     for each color $c$ do
8
           // Block-level parallelization
@@ 3 $\vec{x}\leftarrow$ initial guess with adaptive initialization
                // Thread-level parallelization
@@
11                     $\mathbf{f}_{i,j}=-\frac{\partial E_{j}}{\partial\mathbf{x}_{i}}$
12                     $\mathbf{H}_{i,j}=\frac{\partial^{2}E_{j}}{\partial\mathbf{x}_{i}\partial\mathbf{x}_{i}}$
13
@@ 12                     $\mathbf{H}_{i,j}=\frac{\partial^{2}E_{j}}{\partial\mathbf{x}_{i}\partial\mathbf{x}_{i}}$
15                     $\mathbf{f}_{i}=\sum_{j\in\mathcal{F}_{i}}\mathbf{f}_{i,j}$
16                     $\mathbf{H}_{i}=\sum_{j\in\mathcal{F}_{i}}\mathbf{H}_{i,j}$
17
18                    $\Delta\vec{x}_{i}\leftarrow\mathbf{H}_{i}^{-1}\mathbf{f}_{i}$
19                     $\Delta\vec{x}_{i}\leftarrow$ optional line search from $\vec{x}_{i}+\Delta\vec{x}_{i}$ to $\mathbf{x}_{i}$
20                     $\vec{x}_{i}^{\text{new}}\leftarrow\vec{x}_{i}+\Delta\vec{x}_{i}$
21
@@ 20                     $\vec{x}_{i}^{\text{new}}\leftarrow\vec{x}_{i}+\Delta\vec{x}_{i}$
23
                    // Copy updated positions back to the vertex buffer
@@ 25                          $\mathbf{x}_{i}=\vec{x}_{i}^{\text{new}}$
35                              $\vec{v}=(\vec{x}-\vec{x}^{t})/h$
36                               return $\vec{x}$, $\vec{v}$
Algorithm 1 VBD simulation for one time step.
@@ 35                              $\vec{v}=(\vec{x}-\vec{x}^{t})/h$
                                                                   $\uparrow$ This platform
will be removed.
@@ 5. Results
We use a fixed frame time of $1/60$ seconds and a fixed iteration count $n_{\max}$, instead of estimating convergence after every iteration. Each frame is computed using $S$ substeps, such that ${h=1/(60S)}$ seconds. Using smaller time steps increases accuracy and reduces numerical damping with any implicit Euler method. With VBD, a smaller time step only requires fewer iterations per step for similar visual quality, but it can also achieve a smaller residual with the same number of total iterations per frame (i.e., $Sn_{\max}$). The number of threads per block is set to 16 for all the experiments.
@@ 5.1. Large-Scale Tests
As can be seen in our supplemental video, both of these simulations exhibit stable motion, quickly forming static piles, while maintaining rest-in-contact behavior with over 1 million active collisions. VBD’s parallelism and fast convergence resulted in an average computation time of 40 and 25 seconds per frame in these simulations, respectively.
Figure 13. Visual convergence with different numbers of iterations per frame for different material stiffness (with accelerated iterations using ${\rho=0.75,0.86,0.93}$ top to bottom), simulating a beam with 463 vertices and 1.5 thousand tetrahedra.
@@ 5.4. Convergence Rate
The first alternative is preconditioned gradient descent (GD) (Wang and Yang, 2016) implemented on GPU within the same framework as ours. GD requires a form of line search for stability, which is implemented as testing the variational energy after every 8 iterations and, when needed, reducing the optimization step size and backtracking (following the implementation of Wang and Yang (2016)). We also include a version of GD that is accelerated using the Chebyshev semi-iterative approach (Wang, 2015), as our method. The iterations of GD are about 30% faster than ours without line search. However, it necessitates line search for stability. This makes it about 10% slower than our VBD, which does not need line search. Furthermore, its convergence rate per iteration is considerably slower, as it corresponds to Jacobi iterations. In this example, when combined with acceleration, GD performs similar to our method without acceleration but lags significantly behind our method with acceleration.
@@
We also compare to a version of our method that uses Jacobi iterations, called Block Jacobi, implemented by computing the position change for all vertices in parallel and then applying the position update simultaneously to all vertices at the end of each iteration. We incorporate the same line search scheme as GD for Block Jacobi, as it is necessary for stability. Block Jacobi outperforms GD, as it uses vertex blocks for computation, which corresponds to using a diagonal Hessian block as a precondition, as opposed to just a diagonal line that GD uses. Without line search, it achieves 20% faster iteration times than our VBD, due to its perfect vertex-level parallelization (without any coloring). However, combined with line search, its iterations are about 17% slower than VBD. More importantly, because it uses Jacobi iterations, its convergence is hindered, as compared to our Gauss-Seidel iterations.
@@
Furthermore, we compare the convergence of our method to two implementations of Newton’s method: first using a direct Cholesky (LDLT) factorization solver provided by Intel’s MKL library (Wang et al., 2014) running on the CPU, and the second using a GPU-based conjugate gradient (CG) method. To ensure convergence, we do a PSD projection for each tet’s Hessian of elasticity. Newton’s method uses a line search for each iteration. Since its iterations are slow, the computational overhead of line search is negligible. Though the convergence of Newton’s method per iteration is far superior to all others, because of its expensive computation time per iteration, in these examples it lags far behind. Nonetheless, for relatively tame experiments with less stretching and motion, and especially for highly stiff and high-resolution simulations that are much more expensive to simulate, we would expect Newton’s method to eventually overtake all alternatives beyond a certain level of convergence.
Finally, we compare our method to a quasi-Newton approach using Laplacian preconditioning (Liu et al., 2017). We utilize a GPU-based conjugate gradient solver, similar to the one employed in our GPU-CG Newton’s method’s implementation. Laplacian preconditioning accelerates each linear solve, as it eliminates the need for PSD projection and involves solving a smaller system. Practically, this method demonstrates faster convergence than Newton’s method, particularly in the initial stages of the optimization process. Nevertheless, it still lags behind both the accelerated and non-accelerated versions of our VBD.
Figure 18. Comparing a single-threaded CPU implementation of our method with single-threaded Newton’s method (using both CG and Cholesky). The scene is identical to the bottom row of Figure 17 .
A part of the performance advantages of our VBD method presented above is related to its efficiency in parallel execution on the GPU. To demonstrate its convergence in the absence of parallel computation, we include a comparison using single-threaded CPU implementations of our VBD and Newton’s method with Cholesky factorization and CG. Figure 18 shows the convergence results for the same experiment in the bottom row of Figure 17. In these tests, our method initially demonstrates faster convergence than both versions of Newton’s method. Over time, the CG-based Newton’s method catches up to our VBD without Chebyshev acceleration. However, VBD with Chebyshev acceleration maintains a significant performance lead over the others. This experiment shows that the performance advantages of our method are not only due to its GPU parallelism.
@@
Experiment Name Number of Material Contact &Fiction Simulation Parameters Time per step
@@
Squishy Ball Drops (Figure.7,9,19) 230K 700K 8 NeoHookean $\mu=2e6,\lambda=2e7$ 1e-7 1e7 0.1, 1e-2 1/120 120 15/17ms
@@
2 Cube colliding (Figure 20) 800 2.9K 6 NeoHookean $\mu=1e6,\lambda=1e7$ 1e-6 1e7 0.3, 1e-2 1/300 10 0.16/0/20ms